OpenStudy (anonymous):

if f(x)=x^2 -1. what is the equation for f^-1(x) ....f^-1(x)= +/- (squareroot all)x+1 ...f^-1(x)= +/- (squareroot all )x-1 ....f^-1(x)= +/- (squareroot x only) x+1 ....f^-1(x)= +/- (squareroot x only) x-1

OpenStudy (anonymous):

\[y=x^2-1\] switch x and y \[x=y^2-1\] solve for y \[x+1=y^2\] \[y=\pm\sqrt{x+1}\]

OpenStudy (anonymous):

but this is not a function, so cannot be one of the answers. did they restrict the domain of f?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

then there is a mistake here some place. you cannot write \[f^{-1}(x)=\pm\sqrt{x+1}\] because a function cannot have two answers. on the other hand you are not allowed just to pick one